OpenStudy (anonymous):

Solve by factoring y^2 + 2y = 8

OpenStudy (saifoo.khan):

\[(y+4)(y-2)=0\]y = -4, y =2

OpenStudy (anonymous):

can you show me how you got that?

OpenStudy (saifoo.khan):

Sure. y^2 + 2y = 8 y^2 + 2y - 8 = 0 y^2 +4y -2y -8 = 0 y(y+4)-2(y+4) = 0 (y+4)(y-2) =0 And finally, y = -4,2

OpenStudy (anonymous):

thanks!

OpenStudy (saifoo.khan):

no problem.